# Physics Context Builders: A Modular Framework for Physical Reasoning in Vision-Language Models

### v2

@@ Abstract
Physical reasoning, which involves interpreting object behaviors within dynamic environments, remains a significant challenge for Vision-Language Models (VLMs). The limitations in physical reasoning arise from an inability to translate learned knowledge into predictions about physical behavior. We perform a careful study to show how continual fine-tuning can mitigate this issue. However, fine-tuning is expensive for large models and impractical to repeatedly perform for every task. This necessitates the creation of modular and scalable ways to teach VLMs about physical reasoning. To that end, we introduce Physics Context Builders (PCBs), a novel modular framework where specialized VLMs are fine-tuned to generate detailed physical scene descriptions. These can be used as physical contexts for larger VLMs to enhance their reasoning capabilities. PCBs enable the separation of visual perception from reasoning, allowing us to analyze their relative contributions to physical understanding. We perform careful experiments on CLEVRER and on Falling Tower, a stability detection dataset with both simulated and real-world scenes, to demonstrate that PCBs provide substantial performance improvements, increasing average accuracy by up to 13.8% on complex physical reasoning tasks. Notably, PCBs show strong Sim2Real transfer, successfully generalizing from simulated training data to real-world scenes. Our work demonstrates that enhancing visual perception through modular, simulation-trained components offers a practical approach to improving physical reasoning in VLMs, while providing insights into the factors affecting physical understanding in these models.
@@ 1 Introduction
Physical reasoning is a fundamental component of human intelligence, enabling interpretation of complex interactions, prediction of future events, and understanding of causal relationships in real-world environments (Kubricht et al.,, 2017). For humans, the ability to understand the physical world is developed early and operates intuitively (McCloskey et al.,, 1983; Baillargeon,, 2004; Téglás et al.,, 2011). However, physical reasoning remains a significant challenge for artificial intelligence (AI) systems (Lake et al.,, 2017; Srivastava et al.,, 2022; Yi et al.,, 2019; Chow et al.,, 2025), despite advances in computer vision and Vision-Language Models (VLMs) (Radford et al.,, 2021; Alayrac et al.,, 2022; Li et al.,, 2023; Zhang et al.,, 2024; Sun et al.,, 2024; Li et al.,, 2024; Ventura et al.,, 2025),
@@
VLMs are remarkably successful on many predictive problems, spanning a broad swathe of tasks. This makes their use central to real-world applications requiring expertise in a variety of tasks such as robotics and embodied AI (Liu et al.,, 2024; Guo et al.,, 2024). However, current VLMs consistently fail at physical reasoning tasks, struggling with basic spatial relationships (e.g., object positioning, counting) (Rahmanzadehgervi et al.,, 2024), object attributes (Wang et al.,, 2023), and physical interactions (e.g., stability assessment, dynamics prediction) (Ghaffari and Krishnaswamy,, 2024; Chow et al.,, 2025). While humans leverage causal or physically guided knowledge for physical understanding of the world (Battaglia et al.,, 2013), the mechanisms by which VLMs make predictions are not well understood. This leads to two important questions: what factors contribute to the lack of physical understanding in VLMs and how to improve it?
@@
One potential explanation for such limitations lies in VLMs’ training data. VLMs are created by fusing the representations of image and text encoders using datasets like MSCOCO (Lin et al.,, 2014) and Conceptual Captions (Sharma et al.,, 2018), which focus on general scene descriptions but lack annotations of physical relationships. Our experiments with standard benchmarks like CLEVRER (Yi et al.,, 2019) confirm this hypothesis, demonstrating that fine-tuning on physics-focused data can enable a relatively small VLM to approach state-of-the-art results achieved by specialized architectures (Yi et al.,, 2019; Ding et al., 2021a, ). However, as the ecosystem of open and closed source models grows, repeatedly fine-tuning each VLM for every regime of physical concept it cannot reason about becomes impractical; there is a need for practical, performant and modular tools to augment VLM capabilities.
@@
To this end, we introduce Physics Context Builders (PCBs); these are specialized VLMs that are fine-tuned on simulation data to generate fine-grained physical descriptions that can be used by larger VLMs. Simulation can provide precise annotations of controllable physical interactions of interest, which can serve as scalable training data to teach models about physical phenomena. We show how to amortize the knowledge of physical concepts into PCBs and how they can transfer this knowledge to larger VLMs; this enables us to effectively separate visual perception from reasoning, using PCBs as perception modules, while leaving reasoning to larger models. See Figure 1 for a demonstration of PCBs.
@@
PCBs offer a practical and modular approach to enhance physical reasoning of existing large-scale VLMs without requiring expensive or sometimes infeasible modification of the larger foundation models. Our experiments on standard benchmarks demonstrate effective integration of PCBs with standard commercial VLMs. Moreover, they show that PCBs trained with simulation data successfully generalize to real-world scenarios, effectively performing simulation-to-reality (Sim2Real) transfer. In summary, our contributions are as follows:
@@
    We show how simulation data can be leveraged to train specialized modules that successfully transfer to real-world scenarios, avoiding the need for expensive simulation during inference.
@@ 2 Related Work
Physical Reasoning in Vision Models
@@
Physical reasoning presents significant challenges for vision models, including both specialized architectures and large-scale VLMs. Recent benchmarks demonstrate that current VLMs struggle with basic physics understanding, often failing at tasks such as counting, depth reasoning, and physical interaction prediction (Ghaffari and Krishnaswamy,, 2024; Li et al.,, 2024; Ventura et al.,, 2025). These limitations are not unique to VLMs; physical reasoning has also proven difficult for single-purpose models, necessitating specialized architectures such as physics-inspired predictive models (Battaglia et al.,, 2016; Guen and Thome,, 2020; Duan et al.,, 2022), neural-symbolic executors (Chen et al.,, 2021; Ishay et al.,, 2024; Zheng et al.,, 2024), differentiable physics engines (Wu et al.,, 2015; Ding et al., 2021b, ), simulation-in-the-loop approaches (Liu et al.,, 2022; Zhu et al.,, 2025), and task-specific architectures (Ding et al., 2021a, ). However, large VLMs are not amenable to easy modifications of their neural network architectures and require alternative approaches to enhance their capabilities, which we consider in this work.
Evaluating Physical Reasoning in VLMs
@@
Several recent works have evaluated physical reasoning capabilities in vision-language models. Nagar et al., (2024) benchmark zero-shot visual reasoning in both large language models (LLMs) and VLMs. They find that underlying LLMs, when provided with textual scene descriptions, consistently outperform VLMs that use visual embeddings. Their analysis shows that this performance gap is due to VLMs’ difficulty in translating visual information into accurate representations for reasoning. Our work builds on this insight by developing a modular approach to bridge this gap without requiring extensive retraining of VLMs.
@@
Concurrently, Chow et al., (2025) introduced PhysBench, a comprehensive benchmark for evaluating physical understanding in VLMs. Their work reveals that VLMs’ physical reasoning do not scale proportionally with model size, training data, or input frame count. They identify perceptual and knowledge gaps as the primary sources of errors and propose using vision foundation models like Depth Anything (Yang et al.,, 2024) and SAM (Kirillov et al.,, 2023) to enhance visual perception. While our approach shares the goal of improving physical reasoning, we take a different direction by leveraging simulation data to train specialized context builders rather than relying on generic vision foundation models.
Simulation for Physical Understanding
@@
Simulation has long been recognized as a valuable tool for teaching machines about physical dynamics (Battaglia et al.,, 2013). Previous approaches have incorporated simulation directly into the inference pipeline (Wu et al.,, 2015, 2017; Liu et al.,, 2022), requiring computationally expensive simulators during inference. In contrast, our method leverages simulation only during the training phase to generate rich physical descriptions, eliminating the need for simulation during inference while still benefiting from the detailed annotations that simulations provide.
Enhancing General Capabilities of VLMs
@@
Recent work has explored modular approaches to enhance VLM capabilities without full model retraining. Chain-of-thought prompting techniques (Wei et al.,, 2022) leverage large language models’ reasoning abilities by encouraging step-by-step thinking, while multimodal chain-of-thought approaches (Zhang et al.,, 2023) extend this to vision-language tasks. Our Physics Context Builders build on these ideas by creating modular components that specialize in translating visual inputs into detailed physical descriptions.
@@ 3 Understanding the Effect of Training Data on Physical Reasoning
Here, we investigate why current VLMs struggle with physical reasoning tasks. We first describe our experimental setup, including datasets, models, and training procedures. We then present results demonstrating how fine-tuning with physics-focused data significantly improves performance of VLMs. Finally, we explore the nuanced influence of training data quantity vs. quality.
@@ 3.1 Experimental Setup
We utilize two benchmarks: CLEVRER (Yi et al.,, 2019) for dynamic physical reasoning and Falling Tower for static stability detection.
CLEVRER
is an established benchmark containing 10,000 training videos, 5,000 validation videos, and 5,000 test videos, each paired with multiple questions across four categories: descriptive, explanatory, predictive, and counterfactual. The training set consists of 109,952 descriptive, 16,799 explanatory, 7,179 predictive, and 18,642 counterfactual questions. All descriptive questions are open-ended, while other types are all multi-choice questions with at most four options, where there can be multiple correct options. We report the accuracy of the models stratified by the question types. For multi-choice questions, we report both per-option accuracy, which measures the model’s overall correctness on single options across all questions, and per-question accuracy that measures the model’s performance on full questions, requiring all the choices to be selected correctly. Since the test set answers are not publicly available, we evaluate on the validation set. The dataset’s questions are derived from simulation, making it particularly suitable for our investigation into how simulation-based training data affects physical reasoning.
Falling Tower
Figure 2: Datasets used for experiments.
is a dataset we create to complement CLEVRER by focusing on static physics. Similar to ShapeStacks (Groth et al.,, 2018) and block towers in (Lerer et al.,, 2016), it features stacked objects but extends previous work by including question-answer pairs and simulation-generated annotations. The dataset contains 4,864 images of stacked objects (with 15 object types across 3 shapes and 5 colors) and 72,775 QA pairs, with 75% used for training and 25% for evaluation. Questions and corresponding answers are generated by applying transformation functions to the simulation annotations, converting it into natural language questions and calculating the answers. For instance, to assess the stability of a stacked tower of objects, we compare the initial and final positions of the objects to determine if significant movement occurred. If an object remains within a defined threshold, we conclude it has remained stable. Questions are then categorized into descriptive (e.g., ”How many objects are in the scene?”) and stability (e.g., ”Will this collection of objects remain stationary?”). In addition to the simulated data, the dataset enables the evaluation of Sim2Real transfer; we include 20 real-world images captured with 3D-printed objects with 100 human-generated QAs. An example of the real-world images can be seen in Figure 2. For more details on Falling Tower, see Appendix C.
@@
We evaluate several zero-shot baselines, including GPT-4o, GPT-4o-mini (Achiam et al.,, 2023; Hurst et al.,, 2024), Gemini-1.5-Pro (Team et al.,, 2024), and PaliGemma-3B-mix (a variant fine-tuned on academic datasets) (Beyer et al.,, 2024) with chain-of-thought prompting to encourage explicit reasoning. To study training data effects, we fine-tune PaliGemma-3B on CLEVRER and Falling Tower independently using Low-Rank Adaptation (LoRA) (Hu et al.,, 2021), minimizing the auto-regressive negative log-likelihood of the answers, conditioned on the questions and input video/image. For videos, we sample 8 frames and append it to the input context, except for the Gemini model, where we use the entire video as the input since it supports video input. For all models, we take the final answer after the reasoning chain for evaluation. See section A.1 for more details on the training procedure.
@@
For all questions, including the open-ended and multi-choice ones, we provide the potential options in the question statement. Moreover, for CLEVRER, we frame the multi-choice questions as multiple binary questions, where we ask if each choice can be an answer to the question or not. This yields significant improvement in the accuracy of all the evaluated models, as reported in section B.1.
@@ 3.2 Is Physics-Focused Training Data Enough?
Category Model Descriptive [sim acc. / real acc.] Stability [sim acc. / real acc. / real F1] num. obj. num. $\uparrow\downarrow↑ ↓ obj. absent\uparrow\downarrow$↑ ↓ obj. stable tower stable Random Random 33.3 25.0 14.3 50.0 50.0 Zero-shot CoT GPT-4o 99.3 / 100.0 91.4 / 100.0 99.4 / 95.0 56.9 / 60.0 / 63.3 59.6 / 55.0 / 54.2 GPT-4o-mini 94.9 / 89.5 61.3 / 84.2 87.9 / 73.7 49.0 / 52.6 / 55.0 53.1 / 36.8 / 19.8 Gemini 1.5 Pro 97.2 / 95.0 89.9 / 100.0 97.8 / 95.0 54.6 / 80.0 / 80.0 60.5 / 60.0 / 60.1 PaliGemma-3B-mix 91.4 / 90.0 44.5 / 70.0 73.0 / 65.0 51.0 / 75.0 / 68.6 39.1 / 65.0 / 51.2 Fine-tuned QA PaliGemma-3B-Fine-Tuned 100.0 / 100.0 100.0 / 100.0 100.0 / 95.0 84.6 / 70.0 / 73.0 87.6 / 65.0 / 64.4
@@
Tables 1 and 2 present our results for the Falling Tower and CLEVRER benchmarks, respectively. Several key insights emerge from these results:
@@
On Falling Tower, large models like GPT-4o and Gemini-1.5-Pro demonstrate near-perfect accuracy (95-100%) on descriptive questions, both in simulated and real environments. However, their performance drops substantially on stability questions, with accuracy ranging from only 55-60% (barely above random guessing) on stability detection. Similarly, on CLEVRER, while GPT-4o achieves 62.7% accuracy on descriptive questions, its performance falls to 30.7%, 30.3%, and 18.7% on explanatory, predictive, and counterfactual questions, respectively. This substantial performance gap highlights that while VLMs have developed strong capabilities for high-level scene understanding and description, they lack the abilities needed to predict outcomes or explain causal relationships accurately.
@@
Unsurprisingly, the fine-tuned PaliGemma-3B model shows dramatic improvements across all tasks. On Falling Tower, it achieves perfect descriptive accuracy (100%) and substantial gains in stability prediction (up to 28% improvement compared to GPT-4o). On CLEVRER, the improvements are even more significant, reaching close to the state-of-the-art results by Aloe, a specialized architecture trained on CLEVRER (Descriptive: 94.0 %, Explanatory: 96.0%, Predictive: 87.5%, Counterfactual: 75.6%) (Ding et al., 2021a, ). These results demonstrate that targeted fine-tuning with physics-focused data can significantly enhance physical reasoning capabilities, enabling a relatively small 3B parameter model to outperform much larger state-of-the-art models.

### v3

@@ Abstract
Physical reasoning remains a significant challenge for Vision-Language Models (VLMs). This limitation arises from an inability to translate learned knowledge into predictions about physical behavior. Although continual fine-tuning can mitigate this issue, it is expensive for large models and impractical to perform repeatedly for every task. This necessitates the creation of modular and scalable ways to teach VLMs about physical reasoning. To that end, we introduce Physics Context Builders (PCBs), a modular framework where specialized smaller VLMs are fine-tuned to generate detailed physical scene descriptions. These can be used as physical contexts to enhance the reasoning capabilities of larger VLMs. PCBs enable the separation of visual perception from reasoning, allowing us to analyze their relative contributions to physical understanding. We perform experiments on CLEVRER and on Falling Tower, a stability detection dataset with both simulated and real-world scenes, to demonstrate that PCBs provide substantial performance improvements, increasing average accuracy by up to 13.8% on complex physical reasoning tasks. Notably, PCBs also show strong Sim2Real transfer, successfully generalizing from simulated training data to real-world scenes.
@@ 1 Introduction
Physical reasoning is a fundamental component of human intelligence, enabling interpretation of complex interactions, prediction of future events, and understanding of causal relationships in real-world environments [22]. For humans, the ability to understand the physical world is developed early and operates intuitively [30, 3, 38]. However, physical reasoning remains a significant challenge for artificial intelligence (AI) systems [23, 35, 45, 10], despite advances in computer vision and Vision-Language Models (VLMs) [32, 2, 25, 46, 36, 26, 39].
@@
VLMs are remarkably successful on many predictive problems, spanning a broad range of tasks. This makes their use central to real-world applications requiring expertise in a variety of tasks, such as robotics and embodied AI [28, 17]. However, current VLMs consistently fail at physical reasoning tasks, struggling with basic spatial relationships (e.g., object positioning, counting) [33], object attributes [40], and physical interactions (e.g., stability assessment, dynamics prediction) [14, 10]. While humans leverage causal or physically guided knowledge for physical understanding of the world [5], the mechanisms by which VLMs make predictions are not well understood. This leads to two important questions: what factors contribute to the lack of physical understanding in VLMs, and how can we improve it?
@@
One potential explanation for such limitations lies in VLMs’ training data. VLMs are created by fusing the representations of image and text encoders using datasets like MSCOCO [27] and Conceptual Captions [34], which focus on general scene descriptions but lack annotations of physical relationships. Our experiments with standard benchmarks like CLEVRER [45] confirm this hypothesis, demonstrating that fine-tuning on physics-focused data can enable a relatively small VLM to approach state-of-the-art results achieved by specialized architectures [45, 11]. However, as the ecosystem of open and closed source models grows, repeatedly fine-tuning each VLM for every regime of physical concepts it cannot reason about becomes impractical; there is a need for practical, performant and modular tools to augment VLM capabilities.
@@
To this end, we introduce Physics Context Builders (PCBs): specialized VLMs that are fine-tuned on simulation data to generate fine-grained physical descriptions that can be used by larger VLMs. Simulation can provide precise annotations of controllable physical interactions of interest, which can serve as scalable training data to teach models about physical phenomena. We show how to embed the knowledge of physical concepts into PCBs and how they can transfer this knowledge to larger VLMs; this enables us to effectively separate visual perception from reasoning, using PCBs as perception modules, while leaving reasoning to larger models. See Fig. 1 for a demonstration of PCBs.
@@
PCBs offer a practical and modular approach to enhance the physical reasoning of existing large-scale VLMs without requiring expensive or infeasible modification of the larger foundation models. Our experiments on standard benchmarks demonstrate effective integration of PCBs with standard commercial VLMs. Moreover, they show that PCBs trained with simulation data successfully generalize to real-world scenarios, effectively performing simulation-to-reality (Sim2Real) transfer. In summary, our contributions are:
@@
    We show how simulation can be leveraged to train specialized modules that transfer successfully to real-world scenarios, avoiding expensive simulations at inference.
@@ 2 Related Work
2.1 Physical Reasoning in Vision Models
@@
Physical reasoning presents significant challenges for vision models, including both specialized architectures and large-scale VLMs. Recent benchmarks demonstrate that current VLMs struggle with basic physical understanding, often failing at tasks such as counting, depth reasoning, and physical interaction prediction [14, 26, 39]. These limitations are not unique to VLMs; physical reasoning has also proven difficult for single-purpose models, necessitating specialized architectures such as physics-inspired predictive models [4, 16, 13], neural-symbolic executors [9, 20, 48], differentiable physics engines [42, 12], simulation-in-the-loop approaches [29, 49], and task-specific architectures [11]. However, large VLMs are not easily amenable to architectural modifications and require alternative approaches to enhance their capabilities, which we consider in this work.
2.2 Evaluating Physical Reasoning in VLMs
@@
Several recent works have evaluated physical reasoning capabilities in vision-language models. Nagar et al. [31] benchmark zero-shot visual reasoning in both large language models (LLMs) and VLMs. They find that underlying LLMs, when provided with textual scene descriptions, consistently outperform VLMs that use visual embeddings. Their analysis shows that this performance gap is due to VLMs’ difficulty in translating visual information into accurate representations for reasoning. Our work builds on this insight by developing a modular approach to bridge this gap without requiring extensive retraining of VLMs.
@@
Concurrently, Chow et al. [10] introduce PhysBench, a comprehensive benchmark for evaluating physical understanding in VLMs. Their work reveals that VLMs’ physical reasoning does not scale proportionally with model size, training data, or input frame count. They identify perceptual and knowledge gaps as the primary sources of errors and propose using vision foundation models like Depth Anything [44] and SAM [21] to enhance visual perception. While our approach shares the goal of improving physical reasoning, we take a different direction by leveraging simulation data to train specialized context builders rather than relying on generic vision foundation models.
2.3 Simulation for Physical Understanding
@@
Simulation has long been recognized as a valuable tool for teaching machines about physical dynamics [5]. Previous approaches have incorporated simulation directly into the inference pipeline [42, 43, 29], requiring computationally expensive simulators at inference time. In contrast, our method leverages simulation only during the training phase to generate rich physical descriptions, eliminating the need for simulation during inference while still benefiting from the detailed annotations that simulations provide.
2.4 Enhancing General Capabilities of VLMs
@@
Recent work has explored modular approaches to enhance VLM capabilities without full model retraining. Chain-of-thought prompting techniques [41] leverage large language models’ reasoning abilities by encouraging step-by-step thinking, while multimodal chain-of-thought approaches [47] extend this to vision-language tasks. Our Physics Context Builders build on these ideas by creating modular components that specialize in translating visual inputs into detailed physical descriptions.
@@ 3 Understanding the Effect of Training Data on Physical Reasoning
Figure 2: Datasets used for experiments: Falling Tower representing static physics in simulated and real environments; CLEVRER representing dynamic physics in a simulated environment.
We begin by investigating why current VLMs struggle with physical reasoning. After outlining our experimental setup, covering datasets, models, and training procedures, we present results showing how fine-tuning with physics-focused data significantly improves performance. We then analyze the trade-off between data quantity and quality.
@@ 3.1 Experimental Setup
Datasets.
We utilize two benchmarks: CLEVRER [45] for dynamic physical reasoning and Falling Tower for static stability detection (Fig. 2).
  • CLEVRER [45] is an established benchmark containing 10,000 training videos, 5,000 validation videos, and 5,000 test videos, each paired with multiple questions across four categories: descriptive, explanatory, predictive, and counterfactual. The training set consists of 109,952 descriptive, 16,799 explanatory, 7,179 predictive, and 18,642 counterfactual questions. All descriptive questions are open-ended. Other question types are multiple-choice with up to four options and may have multiple correct answers. We report the accuracy of the models stratified by the question types. For multi-choice questions, we report both per-option accuracy, which measures the model’s overall correctness on single options across all questions, and per-question accuracy that measures the model’s performance on full questions, requiring all the choices to be selected correctly. Since the test set answers are not publicly available, we evaluate on the validation set.
  • Falling Tower is a dataset we create to complement CLEVRER by focusing on static physics. Similar to ShapeStacks [15] and block towers in [24], it features stacked objects but extends previous work by including question-answer pairs and simulation-generated annotations. The dataset contains 4,864 images of stacked objects (with 15 object types across 3 shapes and 5 colors) and 72,775 QA pairs, with 75% used for training and 25% for evaluation. Questions and corresponding answers are generated by applying transformation functions to the simulation annotations, converting it into natural language questions and calculating the answers. For instance, to assess the stability of a stacked tower of objects, we compare the initial and final positions of the objects to determine if significant movement occurred. If an object remains within a predefined threshold, it is considered stable. Questions are then categorized into descriptive (e.g., ”How many objects are in the scene?”) and stability (e.g., ”Will this collection of objects remain stationary?”). This dataset also enables the evaluation of Sim2Real transfer; we include 20 real-world images captured with 3D-printed objects with 100 human-generated QAs. For more details on Falling Tower, see Appendix C.
@@
We evaluate several zero-shot baselines, including GPT-4o, GPT-4o-mini [1, 19], Gemini-1.5-Pro [37], and PaliGemma-3B-mix (a variant fine-tuned on academic datasets) [6] with chain-of-thought prompting to encourage explicit reasoning. To study training data effects, we fine-tune PaliGemma-3B on CLEVRER and Falling Tower independently using Low-Rank Adaptation (LoRA) [18], minimizing the auto-regressive negative log-likelihood of the answers, conditioned on the questions and input video/image. For videos, we sample 8 frames and append them to the input context, except for the Gemini model, where we use the entire video as the input since they support native video processing. For all models, we take the final answer after the reasoning chain for evaluation. Sec. A.1 provides more details on the training procedure.
@@
For all questions, including the open-ended and multi-choice ones, we provide the potential options in the question statement. For CLEVRER, we reframe multi-choice questions as a set of binary questions, asking whether each option is a valid answer. This yields significant improvement in the accuracy of all the evaluated models, as reported in Sec. B.1.
@@ 3.2 Is Physics-Focused Training Data Enough?
Category Model Descriptive [sim acc. / real acc.] Stability [sim acc. / real acc. / real F1] num. obj. num. obj. $\uparrow\downarrow$ obj. $\uparrow\downarrow$ obj. stable tower stable Random Random 33.3 25.0 14.3 50.0 50.0 Zero-shot CoT GPT-4o 99.3 / 100.0 91.4 / 100.0 99.4 / 95.0 56.9 / 60.0 / 63.3 59.6 / 55.0 / 54.2 GPT-4o-mini 94.9 / 89.5 61.3 / 84.2 87.9 / 73.7 49.0 / 52.6 / 55.0 53.1 / 36.8 / 19.8 Gemini 1.5 Pro 97.2 / 95.0 89.9 / 100.0 97.8 / 95.0 54.6 / 80.0 / 80.0 60.5 / 60.0 / 60.1 PaliGemma-3B-mix 91.4 / 90.0 44.5 / 70.0 73.0 / 65.0 51.0 / 75.0 / 68.6 39.1 / 65.0 / 51.2 Fine-tuned QA PaliGemma-3B-Fine-Tuned 100.0 / 100.0 100.0 / 100.0 100.0 / 95.0 84.6 / 70.0 / 73.0 87.6 / 65.0 / 64.4
@@
Tabs. 1 and 2 present our results for the Falling Tower and CLEVRER benchmarks. Several key insights emerge:
@@
On Falling Tower, large models like GPT-4o and Gemini-1.5-Pro demonstrate near-perfect accuracy (95-100%) on descriptive questions, both in simulated and real environments. However, their performance drops substantially on stability questions, with accuracy ranging from only 55-60% (barely above random guessing). Similarly, on CLEVRER, while GPT-4o achieves 62.7% accuracy on descriptive questions, its performance falls to 30.7%, 30.3%, and 18.7% on explanatory, predictive, and counterfactual questions, respectively. This substantial performance gap highlights that while VLMs have developed strong capabilities for high-level scene understanding and description, they lack the abilities needed to predict outcomes or explain causal relationships accurately.
@@
Unsurprisingly, the fine-tuned PaliGemma-3B model shows dramatic improvements across all tasks. On Falling Tower, it achieves perfect descriptive accuracy (100%) and substantial gains in stability prediction (up to 28% compared to GPT-4o). On CLEVRER, the improvements are even more significant, reaching close to the state-of-the-art results by Aloe, a specialized architecture trained on CLEVRER (Descriptive: 94.0 %, Explanatory: 96.0%, Predictive: 87.5%, Counterfactual: 75.6%) [11]. These results demonstrate that targeted fine-tuning with physics-focused data can significantly enhance physical reasoning capabilities, enabling a 3B parameter model to outperform much larger state-of-the-art models. See Sec. D.1 for a comparison to other specialized baselines.